# Puzzle Similarity: A Perceptually-guided Cross-Reference Metric for Artifact Detection in 3D Scene Reconstructions

### v2

@@ Abstract
Modern reconstruction techniques can effectively model complex 3D scenes from sparse 2D views. However, automatically assessing the quality of novel views and identifying artifacts is challenging due to the lack of ground truth images and the limitations of No-Reference image metrics in predicting reliable artifact maps. The absence of such metrics hinders assessment of the quality of novel views and limits the adoption of post-processing techniques, such as inpainting, to enhance reconstruction quality. To tackle this, recent work has established a new category of metrics (Cross-Reference), predicting image quality solely by leveraging context from alternate viewpoint captures [53]. In this work, we propose a new Cross-Reference metric, Puzzle Similarity, which is designed to localize artifacts in novel views. Our approach utilizes image patch statistics from the input views to establish a scene-specific distribution, later used to identify poorly reconstructed regions in the novel views. Given the lack of good measures to evaluate Cross-Reference methods in the context of 3D reconstruction, we collected a novel human-labeled dataset of artifact and distortion maps in unseen reconstructed views. Through this dataset, we demonstrate that our method achieves state of the art localization of artifacts in novel views, correlating with human assessment, even without aligned references. We can leverage our new metric to enhance applications like automatic image restoration, guided acquisition, or 3D reconstruction from sparse inputs. Find the project page at https://nihermann.github.io/puzzlesim/.
@@ 1 Introduction
Image-based rendering and 3D reconstruction from a sparse set of 2D views has received ample attention in recent years, both for pure geometry reconstruction and radiance-field modeling. Classical approaches using simple triangulation and epipolar geometry through methods such as structure from motion (SfM) to produce sparse point-clouds of diffuse color [40]. Densifying these representations can be done explicitly [17]. Alternatively, one can learn continuous, implicit representations [28, 4, 32], normally modeled through some kind of multi-layer perceptron. A tangential problem to these efforts is the collection of 2D data, and the handling of corrupted, distorted, or simply incomplete sets of images from an object or scene we would like to model. Learning representations from very sparse inputs has been a widely studied topic as well [62, 5, 54, 6]; they normally leverage learned priors on large datasets, helping to fill in the gaps by enforcing 3D consistency and that resulting reconstructions follow natural image statistics. However, quantifying the quality of novel views from reconstructions is still problematic. These views can contain artifacts due to the sparsity of the training dataset, and automatically identifying them helps with restoration (e.g. masking for image-based inpainters [44]) or simply to guide future data acquisition to fill the gaps [19]. Recent works have followed a Bayesian approach to quantification of the uncertainty of an area belonging or not to a reconstructed scene [8], which could potentially be leveraged for simple artifact detection; however, they require implicit models, with fundamental changes to the scene model, and are not practical for more general applications that require visual artifact identification outside of scene reconstruction, as well as incapable of detecting artifacts not arising from lack of coverage.
@@
To tackle this, we propose a novel approach for artifact detection that can be leveraged on any set of images without an encoded explicit or implicit model of the scene or object they depict. As opposed to visual difference predictors (VDPs) [23] (which require references) and No-Reference quality metrics [29, 30] (which typically do not provide maps, but rather produce single values of overall quality) our approach provides visual artifact maps with no direct references. We leverage learned perceptual patch statistics from small clean datasets and compare them to the embedded statistics of new images from a similar distribution (i.e. novel reconstructed views from a 3DGS [17] representation) to obtain artifact maps without aligned references. We test our generated maps through a human experiment where we ask participants to manually identify artifacts and distortions in images to generate ground-truth data of visual artifacts. Our results show that our method agrees with human assessment, correlating better than No-Reference, Full-Reference, and state-of-the-art Cross-Reference metrics. To summarize, our contributions are the following:
@@
  • A novel Cross-Reference visual quality/artifact identification metric particularly tailored for 3D reconstruction,
@@
  • a novel dataset of densely human-labeled artifact and distortion maps to fill the gap of validation benchmarks for Cross-Reference metrics,
@@ 2 Related Work
Reconstructing 3D objects or scenes from sparse sets of 2D observations is a fundamental problem in vision [23]. Particularly, in the context of novel view synthesis, the objective is to approximate the radiance field (i.e. 5D function encoding spatially varying radiance emission) of specific objects or scenes. Most methods however, differ either on the model used to encode the function, or the rendering procedure. Implicit approaches model the radiance field as a continuous function, approximated by a multi-layer perceptron [28, 43]. Rendering is usually done via sampling the implicit volume using ray-marching [48], which provides spatially varying values of density and anisotropic color emission modeled through Spherical Harmonics. Improvements over this formula have tackled performance limitations, either by using more efficient sampling techniques [34, 9, 32] or by distilling the implicit space into explicit density and anisotropic appearance volumes [61, 55]. On the other hand, purely explicit models do not require any pre-training using implicit functions, and were originally Eulerian in nature [60]. Explicit models are easier to optimize, usually faster, and more interpretable, which can help in different tasks such as scene editing or animation. More recently, anisotropic Lagrangian approaches have found tremendous success [17]. However, these explicit methods have introduced some limitations of their own along the way. Methods like 3D Gaussian Splatting [17] can only model areas that are directly supervised, and degrade less gracefully than implicit counterparts when querying viewpoints substantially outside the training set coverage. Detecting artifacts arising from the lack of coverage is difficult due to the lack of reference images. Our method produces these masks via supervision on the training data solely, which can enable unsupervised restoration (automatic inpainting of the artifacts based on available context [7]) or simply automatically guide further image acquisition to complete the dataset efficiently [19].
2.2 Image Quality Metrics (IQMs)
Full-Reference Metrics (FR)
FR techniques traditionally input a reference image and its distorted counterpart and compute pixel or patchwise differences, which can either be pooled across an image to provide a single quality value [52, 68] or kept as a pixel-wise map to visualize the location or source of the distortions. Classic examples also include mean-absolute error (MAE), mean-squared error (MSE), or peak signal-to-noise ratio (PSNR). In some cases, it is also possible to transform usually pooled metrics like SSIM [52], FSIM [64], MS-SSIM [51] and LPIPS [68] into pixel maps, and we show how in the Supplementary material. Metrics to evaluate renderings are more commonly yielding maps to localize poorly rendered areas[2].
Visual Difference Predictors (VDPs)
VDPs usually differ from IQMs in the more explicit integration of models of human vision, leveraging the naturally compressive nature of the human visual system [23, 24] to predict perceived differences between images. Improvements over the original framework have increased its robustness by extending their applicability to high dynamic range imagery [26]; making them eccentricity and motion-aware [25]; and integrating perceived color [27]. VDP frameworks have seen use in foveated rendering applications [47] and perceptually-aware tone mappers [46]. In contrast to VDPs, our model relinquishes the necessity of direct references, while performing similarly to state-of-the-art VDPs such as FovVideoVDP [25].
Multi-Modal-Reference Metrics (MMR)
The recent emergence of tasks like image captioning, text-to-image synthesis and text-based image editing demand multimodal quality assessment[39, 15, 3, 35]. Recent MMR metrics leverage joint embedding spaces, such as those learned by CLIP[38], and transformer-based architectures to evaluate the consistency between textual descriptions and visual content[11]. Others localize implausible or misaligned areas of Text-to-Image generated content[22] or assess general image quality[69, 50] from textual templates.
No-Reference Metrics (NR)
No-Reference visual quality metrics relinquish the necessity of direct references and are most commonly learned, often relying on mean-opinion-based supervised learning. Alternatively to leveraging datasets on human quality assessment [16, 14, 58], they can also be supervised on extracted features from natural image statistics [31, 29, 56, 30], or even synthetic scores [57]. Deep learning approaches like TRIQ [59] use Transformer architectures to regress Mean Opinion Scores (MOS) and MUSIQ [16] introduced multi-scale Transformers to alleviate resolution constraints of convolutional networks. Hybrid approaches use multi-modal architectures in conjunction with text templates to query image features like noisiness, sharpness, or contrast that can be translated to MOS[69, 50]. NR metrics that rely on global image features or operate in the frequency domain make it impossible to extract quality maps so we focus on variants that use patch- or pixel-wise statistics.
@@ 2.2 Image Quality Metrics (IQMs)
Localizing artifacts demands IQMs capable of producing visual maps. PIQE [33] attempts to quantify visual artifacts without relying on any supervision; rather, it extracts local features from image patches and quantifies quality as a measure of distortions in the patch, in accordance with a low-level human vision model. CNNIQA [14] was one of the first NR metrics to employ convolutional neural networks (CNN) to regress mean-opinion scores; more recently, PaQ-2-PiQ [58] uses region proposals to select quality-determining patches. PAL4VST [67] localizes specific artifacts that emerge from Image Synthesis Tasks through binary segmentation. In contrast, our work makes no assumptions on the HVS, rather leveraging the latent space of a model pre-trained on natural images to measure the cosine similarity in feature space of candidate image patches to a limited set of image patches from a similar distribution (i.e., images from the same scene in the context of scene reconstruction), rendering a higher level of alignment with human assessment.
Cross-Reference Metrics (CR)
@@
CR methods utilize a set of unaligned reference images to assess the quality of a single query image belonging to the same distribution or ”scene”. They are applicable for novel view synthesis tasks where a scene is generally observed from multiple viewpoints yielding additional information. This category was recently established by Wang et al. [53] and their metric CrossScore serves as the main baseline for our evaluation. CrossScore relies on the cross-attention module[49] to correlate a query image with unaligned multiview images to predict quality maps. The maps predict the quality of 14x14 patches of the input image and are trained to mimic unpooled SSIM maps. However, SSIM has been repeatedly shown to be unaligned with human quality assessment and perception[36, 37, 68], which fundamentally limits the potential of CrossScore. To ensure that our method excels in localizing human apparent artifacts we collected a human-annotated dataset of novel view synthesis artifacts, and demonstrate state-of-the-art results on the task.
@@ 3 Our Method
Choosing the right layers for embedding is essential to maximize the quality of the predicted spatial maps. While early layers feature small receptive fields and capture fine details, deeper layers have larger receptive fields and capture coarser features. This can be observed in Fig. 3, where we showcase different VGG layers. It is essentially a trade-off between prediction granularity, accuracy and speed. We identified that combining multiple layers into our metric computation incorporates the various levels of abstraction and scales in a robust manner. We thus compute the weighted average of the three layers; we empirically found that halving the image resolution more than three times did not significantly improve our results as the scale becomes too small and the pool of reference vectors too little and specific to find good correspondences among novel images, even for well-reconstructed areas. This observation suggested that features from the layers before the third down-sampling were most useful for our cause.
@@
To compute the similarity map of a test image, we feed all references and the test image through a pre-trained network $\mathcal{F}$ to obtain the embeddings. We repeat the exact computation for each network layer, so we will describe the steps once for one layer $\ell$. To find the similarity $s_{\ell}(x,y)$ of the best matching puzzle piece for a pixel of the embedded test image at some spatial location $(x,y)$, we compute the cosine similarity based on the feature vector $\mathcal{F}_{\ell}(x,y)$ and all other feature vectors of all $N$ reference images of the same layer $\ell$ and select the correspondence with the highest similarity:
@@
where $\hat{\mathcal{F}}$ denotes a feature vectors scaled to unit length ^$\hat{\mathcal{F}}_{\ell}(x,y)=\frac{\mathcal{F}_{\ell}(x,y)}{||\mathcal{F}_{\ell}(x,y)||_{2}}\in\mathbb{R}^{C_{\ell}}$ and $\cdot$ is the dot product. Note that we compute the cosine similarity with any feature vector of the same layer from all references, not just those at the same spatial position. This relinquishes spatial relations and makes the method robust to simple camera movements that only shift the image horizontally or vertically. We iterate this maximum search for all pixels of the test image’s embedding to construct the similarity mask $S_{\ell}$.
$S_{\ell}(\mathcal{I})=\begin{bmatrix}s_{\ell}(1,1)&s_{\ell}(1,2)&\cdots&s_{\ell}(1,W_{\ell})\\ s_{\ell}(2,1)&s_{\ell}(2,2)&\cdots&s_{\ell}(2,W_{\ell})\\ \vdots&\vdots&\ddots&\vdots\\ s_{\ell}(H_{\ell},1)&s_{\ell}(H_{\ell},2)&\cdots&s_{\ell}(H_{\ell},W_{\ell})\\ \end{bmatrix}$ (2)
where $\mathcal{I}$ is the test image. We repeat this for a set of layers and combine them into a final quality map. To match the spatial dimensions of each layer, we bilinearly upsample each map to the original image size and combine them with an affine combination:
$S(\mathcal{I})$ $=\sum_{\ell}w_{\ell}\;\text{Upsample}\left(S_{\ell}(\mathcal{I},\;\mathcal{I}_{\text{ref}}^{1:N})\right)$ (3)
@@
The choice of pre-trained neural network, through which the embeddings will be created, is a key component of our work. We primarily considered classic models including VGG-16, VGG-19, AlexNet, and SqueezeNet [42, 20, 12]. Some of the critical considerations are model complexity and memory requirements, which we summarized in the Supplemental, as well as their specific tested performance in our human assessment alignment task. Beyond quality performance, reducing the memory footprint and computational complexity is key as it may impact the possibility of downstream applications of our metric, which, given its differentiability, could be leveraged in optimization procedures.
We empirically found that while VGG produces the most fine-grained maps, AlexNet and SqueezeNet still managed to perform similarly; while doing so at a substantially reduced computational cost. We opted for SqueezeNet as it aligned best with our test examples, specifically using layers $\ell\in\{2,3,4\}$ with the weights $w_{2}=0.67$, $w_{3}=0.2$, and $w_{4}=0.13$, which we found heuristically.

### v3

@@ Abstract
Modern reconstruction techniques can effectively model complex 3D scenes from sparse 2D views. However, automatically assessing the quality of novel views and identifying artifacts is challenging due to the lack of ground truth images and the limitations of no-reference image metrics in predicting reliable artifact maps. The absence of such metrics hinders assessment of the quality of novel views and limits the adoption of post-processing techniques, such as inpainting, to enhance reconstruction quality. To tackle this, recent work has established a new category of metrics (cross-reference), predicting image quality solely by leveraging context from alternate viewpoint captures [47]. In this work, we propose a new cross-reference metric, Puzzle Similarity, which is designed to localize artifacts in novel views. Our approach utilizes image patch statistics from the training views to establish a scene-specific distribution, later used to identify poorly reconstructed regions in the novel views. Given the lack of good measures to evaluate cross-reference methods in the context of 3D reconstruction, we collected a novel human-labeled dataset of artifact and distortion maps in unseen reconstructed views. Through this dataset, we demonstrate that our method achieves state-of-the-art localization of artifacts in novel views, correlating with human assessment, even without aligned references. We can leverage our new metric to enhance applications like automatic image restoration, guided acquisition, or 3D reconstruction from sparse inputs. Find the project page at https://nihermann.github.io/puzzlesim/.
@@ 1 Introduction
Image-based rendering and 3D reconstruction from a sparse set of 2D views has received ample attention in recent years, both for pure geometry reconstruction and radiance-field modeling. Classical approaches such as structure from motion (SfM) use simple triangulation and epipolar geometry to produce sparse point clouds of diffuse color [33]. Densifying these representations can be done explicitly [15]. Alternatively, one can learn continuous, implicit representations [24, 3, 28], normally modeled through multi-layer perceptrons. A tangential problem to these efforts is the collection of 2D data and the handling of corrupted, distorted, or simply incomplete sets of images from an object or scene we would like to model. Learning representations from very sparse inputs has been a widely studied topic, where normally learned priors from large datasets are leveraged to enforce 3D consistency to ensure that the resulting reconstructions follow natural statistics [57, 4, 48, 5]. However, quantifying the quality of novel views from reconstructions is still problematic. These views can contain artifacts due to the sparsity of the training dataset, and automatically identifying them helps with restoration (e.g., masking for image-based inpainters [37]) or simply to guide future data acquisition to fill the gaps [17]. Recent works have followed a Bayesian approach to quantify the uncertainty of whether an area belongs to a reconstructed scene or not [8], which could potentially be leveraged for simple artifact detection. However, they require implicit models, with fundamental changes to the scene model, and are not practical for more general applications that require visual artifact identification outside of scene reconstruction, and are incapable of detecting artifacts not arising from lack of coverage.
@@
To tackle this, we propose a novel approach for artifact detection that can be leveraged on any set of images without an encoded explicit or implicit model of the scene or object they depict. Unlike visual difference predictors (VDPs) [20] (which require references) and no-reference quality metrics [25, 26] (which typically do not provide maps, but rather produce single values of overall quality), our approach provides visual artifact maps with no direct references. We leverage learned perceptual patch statistics from small, clean datasets and compare them to the embedded statistics of new images from a similar distribution (i.e., novel reconstructed views from a 3DGS [15] representation) to obtain artifact maps without aligned references. We test our generated maps through a human experiment where we ask participants to manually identify artifacts and distortions in images to generate ground-truth data of visual artifacts. Our results show that our method agrees with human assessment, correlating better than no-reference, full-reference, and state-of-the-art cross-reference metrics. To summarize, our contributions are the following:
@@
  • A novel cross-reference visual artifact identification metric, particularly tailored for 3D reconstruction,
@@
  • a novel dataset of human-labeled artifact and distortion maps to fill the gap of validation benchmarks for cross-reference metrics,
@@ 2 Related Work
Our metric is specifically designed for applications in 3D scene reconstruction and image-based rendering. Consequently, this section discusses work on 3D reconstruction first and then on image metrics.
@@
Reconstructing 3D objects or scenes from sparse sets of 2D observations is a fundamental problem in vision [20]. Particularly, in the context of novel view synthesis, the objective is to approximate the radiance field (i.e., 5D function encoding spatially varying radiance emission) of specific objects or scenes. Most methods differ either in the model used to encode the function or the rendering procedure. Implicit approaches model the radiance field as a continuous function, approximated by a multi-layer perceptron [24, 36]. Rendering is usually done via sampling the implicit volume using ray-marching [42], which provides spatially varying values of density and anisotropic color emission modeled through Spherical Harmonics. Improvements over this formula have tackled performance limitations, either by using more efficient sampling techniques [30, 9, 28] or by distilling the implicit space into explicit density and anisotropic appearance volumes [56, 50]. On the other hand, purely explicit models do not require any pre-training using implicit functions, and were originally Eulerian in nature [55]. Explicit models are easier to optimize, usually faster, and more interpretable, which can help in different tasks such as scene editing or animation. More recently, anisotropic Lagrangian approaches have found tremendous success [15]. However, these explicit methods have introduced some limitations of their own along the way. Methods like 3D Gaussian Splatting [15] can only model areas that are directly supervised, and degrade less gracefully than implicit counterparts when querying viewpoints substantially outside the training set coverage. Detecting artifacts arising from the lack of coverage is difficult due to the lack of reference images. Our method produces these masks via supervision on the training data solely, which can enable unsupervised restoration (automatic inpainting of the artifacts based on available context [7]) or simply automatically guide further image acquisition to complete the dataset efficiently [17].
2.2 Image Metrics
Image metrics are classified by the type of prediction they make and their input. Image Quality Metrics (IQMs) typically predict a single number, corresponding to overall image quality, and are often trained on Mean Opinion Score (MOS) datasets. Visibility Metrics (VMs) produce maps corresponding to the perceptibility of distortions. They often rely on models of the human visual system and predict the probability of detecting local artifacts by an observer. Most image metrics are full-reference, necessitating a reference to assess the quality of a test image. In contrast, no-reference metrics predict the quality or visible distortions based solely on the test image. Others use additional information, e.g., partial reference, and are referred to as reduced-reference or cross-reference metrics.
Classical examples of full-reference IQMs include mean-absolute error (MAE), mean-squared error (MSE), peak signal-to-noise ratio (PSNR), SSIM [46], FSIM [59], MS-SSIM [45], and LPIPS [62]. These methods begin by computing local differences between test and reference images and aggregate them into a single quality score as the final step. By omitting this step, it is possible to create a local distortion map, which is a common output of VMs widely adopted in rendering to localize poorly rendered areas [2]. Typically, VMs differ from IQMs in their more explicit modeling of the human visual system [6], which enables the prediction of perceived visibility or the strength of visual differences between images. Improvements over the original framework extended their applicability to high dynamic range imagery [22], making them eccentricity and motion-aware [40, 21], and integrating perceived color [23]. Such metrics have been used in many perceptual optimizations, such as foveated rendering [41] and perceptually aware tone mappers [39].
@@ 2.2 Image Metrics
No-reference metrics eliminate the need for a reference and are most commonly learned from human quality assessment datasets [14, 13, 53], supervised on extracted features from natural image statistics [27, 25, 51, 26], or even synthetic scores [52]. Modern deep learning approaches can utilize Transformer architectures [54], and multi-scale Transformers are employed to alleviate the resolution constraints [14]. Hybrid approaches use multi-modal architectures in conjunction with text templates to query image features such as noisiness, sharpness, or contrast, which can be translated into MOS [63, 44]. The above no-reference metrics rely on global image features and, therefore, are not suitable for obtaining distortion maps. However, similarly to our work, some of these IQMs are capable of producing visual maps. For example, PIQE [29] measures distortions in an image patch based on extracted local features. CNNIQA [13], on the other hand, was one of the first no-reference metrics to employ convolutional neural networks (CNN) to regress mean-opinion scores. More recently, PaQ-2-PiQ [53] uses region proposals to select quality-determining patches. PAL4VST [61] localizes specific artifacts that emerge from image synthesis tasks through binary segmentation. In contrast, our work leverages the latent space of a model pre-trained on natural images to measure the cosine similarity in feature space of candidate image patches to a limited set of image patches from a similar distribution (i.e., images from the same scene in the context of scene reconstruction), rendering a higher level of alignment with human assessment.
@@
In some applications, even though an image metric does not have access to a reference image, it may have access to other information useful for making a prediction. In the context of novel view synthesis, cross-reference metrics utilize a set of unaligned reference images to assess the quality of a single query image belonging to the same scene. This type of metric was recently established by Wang et al. [47] and their metric CrossScore serves as the main baseline for our evaluation. The metric relies on a cross-attention module [43] to correlate a test image with unaligned multiview images to predict quality maps. The maps predict the quality of 14x14 patches of the input image and are trained to mimic unpooled SSIM maps. However, SSIM has been repeatedly shown to be poorly aligned with human quality assessment and perception [31, 32, 62], which fundamentally limits the potential of CrossScore.
In comparison to previous work, our method is a cross-reference metric for novel view synthesis tasks. It predicts local similarities for a synthesized view given a set of unaligned reference views. Its effectiveness in localizing artifacts is achieved by assessing similarity in feature space. In contrast to many IQMs, our method is not designed to predict a single score, but rather a map corresponding to the strength of the visible artifacts.
@@ 3 Our Method
Choosing the right layers for embedding is essential to maximize the quality of the predicted spatial maps. While early layers feature small receptive fields and capture fine details, deeper layers have larger receptive fields and capture coarser features. This can be observed in Fig. 3, where we showcase different VGG layers. It is essentially a trade-off between prediction granularity, accuracy and speed. We identified that combining multiple layers into our metric computation incorporates the various levels of abstraction and scales in a robust manner. We thus compute the weighted average of the three layers; we empirically found that halving the image resolution more than three times did not significantly improve our results, as the scale becomes too small and the pool of reference vectors too little and specific to find good correspondences among novel images, even for well-reconstructed areas. This observation suggested that features from the layers before the third down-sampling were most useful for our cause.
@@
[…]
@@
where $\hat{\mathcal{F}}$ denotes a feature vectors scaled to unit length ^$\hat{\mathcal{F}}_{\ell}(x,y)=\frac{\mathcal{F}_{\ell}(x,y)}{||\mathcal{F}_{\ell}(x,y)||_{2}}\in\mathbb{R}^{C_{\ell}}$ and $\cdot$ is the dot product. Note that we compute the cosine similarity with any feature vector of the same layer from all references, not just those at the same spatial position. This relinquishes spatial relations and makes the method robust to simple camera movements that only shift the image horizontally or vertically. We iterate this maximum search for all pixels of the test image’s embedding to construct the similarity mask $\mathcal{S}_{\ell}$.
$\mathcal{S}_{\ell}(\mathcal{I})=\begin{bmatrix}s_{\ell}(1,1)&s_{\ell}(1,2)&\cdots&s_{\ell}(1,W_{\ell})\\ s_{\ell}(2,1)&s_{\ell}(2,2)&\cdots&s_{\ell}(2,W_{\ell})\\ \vdots&\vdots&\ddots&\vdots\\ s_{\ell}(H_{\ell},1)&s_{\ell}(H_{\ell},2)&\cdots&s_{\ell}(H_{\ell},W_{\ell})\\ \end{bmatrix}$ (2)
where $\mathcal{I}$ is the test image. We repeat this for a set of layers and combine them into a final similarity map. To match the spatial dimensions of each layer, we bilinearly upsample each map to the original image size and combine them with an affine combination:
$\mathcal{S}(\mathcal{I})$ $=\sum_{\ell}w_{\ell}\;\text{Upsample}\left(\mathcal{S}_{\ell}(\mathcal{I},\;\mathcal{I}_{\text{ref}}^{1:N})\right)$ (3)
@@
The choice of pre-trained neural network, through which the embeddings will be created, is a key component of our work. We primarily considered classic models including VGG-16, VGG-19, AlexNet, and SqueezeNet [35, 18, 11]. Some of the critical considerations are model complexity and memory requirements, which we summarized in the Supplemental, as well as their specifically tested performance on our human alignment task. Beyond quality performance, reducing the memory footprint and computational complexity is key as it may impact the possibility of downstream applications of our metric, which, given its differentiability, could be leveraged in optimization procedures.
We empirically found that while VGG produces the most fine-grained maps, AlexNet and SqueezeNet still managed to perform similarly, while doing so at a substantially reduced computational cost. We opted for SqueezeNet as it aligned best with our test examples, specifically using layers $\ell\in\{2,3,4\}$ with the weights $w_{2}=0.67$, $w_{3}=0.2$, and $w_{4}=0.13$, which we found heuristically.